Shirel burns 10 mg of phosphorus in the presence of 1.5 mg of oxygen gas in a closed beaker. This reaction produces diphosphorus pentoxide. What is the correct chemical formula for the chemical produced in the reaction?


1
Expert's answer
2021-03-25T04:11:11-0400

P4 + O2 -> P2O5

10 mg of phosphorus

Molar Mass of Phosphorus = 30.97

10/30.97 = 1/3 P = 1/12 P4

1.5 mg of Oxygen

Molar Mass of Oxygen = 15.999

1.5/15.999 = 5/33 O = 1/16 O2

= 1/12 P4 + 1/16 O2 -> 1/4P2O5

= 1/3P4 + 1/4O2 -> P2O5